OpenStudy (rebecca1588):

***FAN AND MEDAL*** Which is an equation in point-slope form for the given point and slope? Point: (8, -4) ; Slope: -5/6 A. y+4=-5/6(x+8) B. y-4=-5/6(x+8) C .y+4=-5/6(x-8) D. y-4=5/6(x-8)

OpenStudy (dan815):

hi

OpenStudy (rebecca1588):

Haii

OpenStudy (dan815):

y-b=x-a , is the line y=x shifted to the point (a,b)

OpenStudy (dan815):

y-b=m*(x-a) is the line y=mx, with slope m shifted to point (a,b)

OpenStudy (rebecca1588):

It's C right?

OpenStudy (dan815):

yes

OpenStudy (rebecca1588):

Definitely C. You just have to plug in the numbers in the point slope formula with 8 as x sub one, -4 as y sub one, and -5/6 as your m or slope.

OpenStudy (dan815):

yep

OpenStudy (dan815):

y-b=m*(x-a) is the line y=mx, with slope m shifted to point (a,b) m=-5/6, (a,b)=(8,-4) subbing this in y-(-4) = -5/6 (x-8) y+4=-5/6(x-8)

OpenStudy (rebecca1588):

Thanks for your help! ^-^

OpenStudy (dan815):

you're welcome! ^~^